OpenStudy (anonymous):

can some one Find the next three terms in the sequence plz 3, 12, 21, 30, . . .

OpenStudy (anonymous):

\[a_n = 3(3n-2)\] \[n = term~you~want\] You know that the fourth term is 30 you want the 5, 6, 7 term then. \[a_5 = 3(3(5)-2) \rightarrow 39\] \[a_6 = 3(3(6)-2) \rightarrow 48\] \[a_7 = 3(3(5)-2) \rightarrow 57\]